# Dexterous Manipulation through Imitation Learning: A Survey

### v3

@@ IV End Effectors for Dexterous Manipulation
An end effector is a component at the tip of a robotic manipulator that directly interacts with the environment to perform tasks. In dexterous manipulation, end effectors are typically categorized into two-fingered grippers, multi-fingered anthropomorphic hands, and three-fingered robotic claws. This section introduces these three types in order, highlighting their design principles, advantages, and trade-offs.
@@ IV-B Multi-fingered Anthropomorphic Hand
Figure 3: Examples of multi-fingered anthropomorphic hands: (a) Shadow Dexterous Hand[119]; (b) Awiwi Hand [120]; (c) Biomimetic Hand [121]; (d) ILDA Hand [117]; (e) Hu et al.’s robotic hand [118]; (f) INSPIRE-ROBOTS RH56 Dexterous Hand[122]; (g) OYMotion OHand[123]; (h) PUT-Hand [124]; (i) Allegro Hand[125]; (j) Faive Hand [126]; (k) Tesla Optimus Hand[127]; (l) Utah/MIT Dexterous Hand [128].
@@ IV-B2 Linkage-driven Approach
In the one-DoF type, joints are mechanically coupled, so preshaping remains fixed during flexion. Typical designs include the S-finger with inverse four-bar coupling [163] and the humanoid hand by Liu et al. using two four-bar linkages per finger [164]. Similar configurations appear in hands like the INSPIRE-ROBOTS RH56 [122] (see Fig. 3(f)), Bebionic Hand [165, 166], BrainRobotics Hand [167] and OYMotion OHand [123] (see Fig. 3(g)).
@@
While most designs use one motor per finger, a few incorporate multiple actuators for higher dexterity. The ILDA Hand [117] (see Fig. 3(d)) employs three motors per finger with combined PSS/PSU chains and four-bar linkages, achieving workspace and fingertip force comparable to human hands. Similar high-DoF linkage designs appear in the AIDIN ROBOTICS Hand [173] and the RY-H1 Hand [174].
@@ IV-C Three-fingered Robotic Claw: A Trade-off Solution
Figure 4: Examples of three-fingered robotic claws: (a) DEX-EE [204]; (b) BarrettHand [205]; (c) i-HY Hand [199]; (d) DoraHand [206].
@@
Three-fingered claws such as the DoraHand [203] (see Fig. 4(d)), SARAH [215], D’Manus [216], and Kinova Jaco’s claw [217] further demonstrate the practicality and versatility of this design choice in both research and assistive applications.
V Teleoperation Systems and Data Collection
@@ V-A Teleoperation Systems for Dexterous Manipulation
To accurately convey human operators’ intentions to robotic systems, previous works have employed a wide range of human-robot interaction devices. Human operators with work experience can easily identify the current state of a robot through the image; however, accurately translating human instructions to robot actions remains a challenge. Some traditional controllers act on this: 1) joysticks [218] 2) haptic devices [219]; However, manipulation tasks often involve delicate movements and complex interactions, such as grasping, moving, and positioning small or irregularly shaped objects. These tasks necessitate devices that can offer dexterous interfaces to ensure the safety and efficacy of the robot’s actions. Precision and real-time feedback are crucial. Commonly used devices include: 1) cameras [18, 220, 17, 19, 221, 222, 223, 54, 20]; 2) mocap gloves [224, 225, 226, 227, 228, 229]; 3) VR/AR controllers [230, 231, 14, 232, 233, 54, 234, 27, 235, 236, 237, 238]; 4) exoskeletons and bilateral systems [239, 240, 241, 53, 242, 243].
@@ V-A4 Exoskeleton and Bilateral Systems
V-B Datasets and Benchmarks
V-B1 Datasets
MIME [250]is a large-scale dataset containing 8,260 human-robot demonstrations across 20 diverse tasks ranging from simple tasks like pouring to difficult tasks like stacking objects. It includes both videos of human demonstrations and kinesthetic trajectories for robots.
RH20T [251] dataset encompasses over 110,000 multimodal robotic manipulation sequences collected using intuitive teleoperation interfaces equipped with force-torque sensors and haptic feedback. It captures visual, tactile, audio, and proprioceptive data alongside human demonstration videos, facilitating one-shot IL across diverse tasks and robotic configurations. This dataset aims to advance robotic skill acquisition in unstructured environments through enhanced task and motion planning.
BridageData [252] encompasses 7,200 demonstrations across 71 tasks in 10 environments. Leveraging diverse and complex manipulations primarily in kitchen settings. It aims to support broad skill generalization via cross-domain datasets. Following this, BridgeData V2 [253] expands the range of tasks and environments to foster more robust generalization and transfer capabilities in robots. With 60,096 robotic manipulation trajectories across 24 environments, it supports scalable robot learning with tasks varying from pick-and-place to complex manipulation, facilitating generalization across tasks, objects, and settings for multi-task and language-conditioned learning methods.
DROID [254] surpasses other datasets collected purely via human teleoperation with its unparalleled scene diversity and task variety. It is a diverse robot imitation learning dataset with 76,000 demonstration trajectories across 86 tasks and 564 scenes.
To address the problem that collecting extensive human demonstrations for IL is always time-consuming and laborious, several datasets apply data augmentation to demonstrations collected via human teleoperation. To illustrate, data augmentation is a technique that has been widely applied in various fields, particularly in computer vision and robotics. It is used to artificially expand the size and diversity of a training dataset by applying a series of transformations or modifications to the original data, which can improve the robustness and generalization capabilities of machine learning models trained on the augmented dataset.
RoboAgent [255] contains 7,500 trajectories collected through human teleoperation. It can be scaled up to approximately 98,000 trajectories to diversify data through semantic augmentations without extra human/robot cost. Similarly, CyberDemo [256] follows this approach. Researchers collect human demonstrations using teleoperation in both simulated and real-world environments and then implement extensive data augmentation to the demonstrations collected. By incorporating visual and physical variations in simulations, robots trained in them could obtain enhanced policy robustness and generalization ability.
Several other datasets leverage the demonstration generation system to enhance IL by expanding the dataset from a limited number of human demonstrations. In MimicGen [257], over 50,000 demonstrations are created across 18 tasks from approximately 200 human demonstrations; it synthesizes full demonstrations of diverse scene configurations, object instances, and robot arms by adapting object-centric manipulation behaviors to new contexts through trajectory transformations based on known object poses, enabling the training for complex, long-horizon, and high-precision tasks.
Similarly, IntervenGen[258], introduced in 2024, also autonomously produces large sets of corrective interventions from minimal human input, enhancing policy robustness against distribution shifts. It specializes in generating interventional data to address policy mistakes, further reducing human effort and improving robustness.
DiffGen [259] basically follows the very same idea but integrates differentiable physics simulation, rendering, and vision-language models to generate realistic robot demonstrations from text-based instructions.
There are also datasets that focus on dexterous bimanual manipulations and hand-object interactions. For example, ARCTIC [260] contains 2.1 million videos with accurate 3D hand-object meshes and dynamic contact data, enabling the study of dexterous bimanual manipulation of articulated objects. It introduces novel tasks for consistent motion reconstruction and interaction field estimation, facilitating advanced research in hand-object interaction.
Additionally, DexGraspNet [261] featuring 1.32 million grasps of 5,355 objects by ShadowHand fills the blank of large-scale, diverse, and high-quality datasets for dexterous grasping. It includes over 200 diverse grasps per object, validated for physical stability in simulation, enabling more effective IL and benchmarking of robotic manipulation algorithms to achieve human-like dexterity and grasping capabilities.
The OAKINK2 dataset [262] comprises 627 sequences of bimanual object manipulation, featuring 4.01 million frames from multi-view captures and detailed pose annotations for human bodies, hands, and objects.

### v4

@@ IV End-effectors for Dexterous Manipulation
An end-effector is a component at the tip of a robotic manipulator that directly interacts with the environment to perform tasks. In dexterous manipulation, end-effectors are typically categorized into two-fingered grippers, multi-fingered anthropomorphic hands, and three-fingered robotic claws. This section introduces these three types in order, highlighting their design principles, advantages, and trade-offs. Table IV provides a comparative analysis of the performance characteristics of these three end-effector types in dexterous manipulation.
@@ IV-B Multi-fingered Anthropomorphic Hand
Figure 3: Examples of multi-fingered anthropomorphic hands: (a) Shadow Dexterous Hand[150]; (b) Awiwi Hand [151]; (c) Biomimetic Hand [152]; (d) ILDA Hand [148]; (e) Hu et al.’s robotic hand [149]; (f) INSPIRE-ROBOTS RH56 Dexterous Hand[153]; (g) Linker Hand L20 [154]; (h) PUT-Hand [155]; (i) Allegro Hand[156]; (j) Faive Hand [157]; (k) Tesla Optimus Hand[158]; (l) Utah/MIT Dexterous Hand [159].
@@ IV-B2 Linkage-driven Approach
In the one-DoF type, joints are mechanically coupled, so preshaping remains fixed during flexion. Typical designs include the S-finger with inverse four-bar coupling [193] and the humanoid hand by Liu et al. using two four-bar linkages per finger [194]. Similar configurations appear in hands like the INSPIRE-ROBOTS RH56 [153] (see Fig. 3(f)), Bebionic Hand [195, 196], BrainRobotics Hand [197] and OYMotion OHand [198].
@@
While most designs use one motor per finger, a few incorporate multiple actuators for higher dexterity. The ILDA Hand [148] (see Fig. 3(d)) employs three motors per finger with combined PSS/PSU chains and four-bar linkages, achieving workspace and fingertip force comparable to human hands. Similar high-DoF linkage designs appear in the Linker Hand L20 [154](see Fig. 3(g)), AIDIN ROBOTICS Hand [204] and the RY-H1 Hand [205].
@@ IV-B4 Hybrid-transmission Approach
TABLE V: Key Characteristics of Representative Robotic Hands
Robotic Hand DoF (actuated / total) Actuation Type Key Feature Typical Applications
Shadow Dexterous Hand[150] 20 / 24 Remote tendon-driven (20 Smart Motors in forearm) Fully biomimetic; 129 sensors; $\pm$1° joint precision High-dexterity research, tele-operation, AI/ML benchmarking
Allegro Hand[156] 16 / 16 Direct-driven (Maxon DC motors in each phalanx) Modular finger units; ROS-native; low backlash Academic RL/IL experiments, sim-to-real transfer
Linker Hand (L20)[154] 16 / 21 Linkage-driven (4 motors per finger) ROS-native; piezoresistive sensors; optional visual-tactile perception High-dexterity research, Industrial applications
BarrettHand (BH8-280)[222] 4 / 7 Linkage-driven (1 motor per finger + spread) Under-actuated adaptive fingers; rugged mechanics Industrial grasping, pick-and-place, educational labs
LEAP Hand[209] 16 / 16 Direct-driven (compact joint motors) Low-cost, open-source, anthropomorphic yet lightweight Learning-focused projects, mobile manipulation
i-HY Hand (3-finger)[223] 9 / 9 Tendon-driven (elastic tendons + capstan) Compliant, under-actuated; high-impact robustness Robust in-hand manipulation, field robotics
@@ IV-C Three-fingered Robotic Claw: A Trade-off Solution
Figure 4: Examples of three-fingered robotic claws: (a) DEX-$E$ [236]; (b) BarrettHand [222]; (c) i-HY Hand [223]; (d) DoraHand [237].
@@
Three-fingered claws such as the DoraHand [235] (see Fig. 4(d)), SARAH [246], D’Manus [247], and Kinova Jaco’s claw [248] further demonstrate the practicality and versatility of this design choice in both research and assistive applications. Based on the foregoing analysis, the key characteristics of representative robotic hands across the three end-effector categories are summarized in Table V.
IV-D Tactile Sensors on Dexterous End-effectors
Tactile sensing is important for enhancing the perception and control capabilities of robotic end-effectors, particularly in dexterous manipulation tasks that require fine contact interaction, compliance, and force regulation. Although previous end-effectors such as two-finger grippers and parallel jaw mechanisms often rely solely on vision or position feedback, modern multi-fingered dexterous hands [249, 250] are increasingly incorporating high-resolution tactile sensors to achieve more stable and versatile manipulation in unstructured or dynamic environments.
In general, tactile sensors provide direct measurements of contact force, pressure distribution, slip, surface texture, and sometimes even temperature or material properties. They are typically mounted on the fingertips, phalanges, or palm surfaces of robotic hands, allowing feedback-rich control policies such as force modulation, slip prevention, and adaptive grasping. In recent years, tactile sensors have become essential for learning-based approaches, where tactile data are used for contact state estimation, affordance recognition, and policy refinement [251, 252, 253, 254].
Tactile sensors can be broadly categorized into several types based on sensing principles:
  • Resistive and Capacitive Tactile Sensors: These are thin, flexible, and cost-effective, and detect pressure changes through variations in electrical resistance or capacitance. They are widely used in fingertip arrays and robotic skin [255, 256, 257].
  • Piezoelectric Sensors: These generate voltage in response to applied stress and are suitable for dynamic contact detection, such as impact and slip sensing [258, 259].
  • Optical and Vision-Based Tactile Sensors: Examples such as GelSight [260], TacTip [261], and GelStereo tip [262] use internal cameras and deformable gel surfaces to reconstruct high-resolution 3D contact geometry. These sensors offer rich, interpretable data and are increasingly used in learning-based manipulation [263].
  • Magnetic and Hall Effect Sensors: Often embedded within soft materials or joints, these detect local deformations or forces via magnetic field changes [264, 265, 266, 267].
  • Bio-inspired Sensors: Inspired by human skin or mechanoreceptors, these sensors aim to capture multi-modal tactile features (e.g., pressure + vibration + shear) in compact designs. Examples include BioTac [268], NeuTouch [269], and GTac [270].
The integration of tactile sensing into robotic hands offers significant advantages. First, it enables robust manipulation under conditions of visual occlusion or low illumination by relying solely on contact feedback. Second, tactile sensors facilitate adaptive grasping strategies such as regrasping, force-limited lifting, and slip compensation. Third, tactile data contribute to improved grasp quality estimation, object property inference, and planning for in-hand manipulation. However, tactile sensing also faces several challenges. Calibration and drift over time, especially in soft or flexible sensors, can reduce measurement fidelity. High-resolution sensors (e.g. GelSight [260]) are relatively bulky and have limited durability under repeated contact. Furthermore, real-time integration of tactile data into control pipelines demands efficient signal processing, accurate contact modeling, and often machine learning-based interpretation.
In the context of IL for dexterous manipulation, tactile sensing provides a rich stream of contact information that complements visual and proprioceptive cues. Demonstrations collected from expert teleoperation or kinesthetic teaching can include synchronized tactile signals, allowing the learning agent to capture subtle contact events, such as incipient slip, rolling, or micro-adjustments of finger force, that are often invisible to vision alone. By incorporating tactile feedback into the policy representation, IL systems can better generalize to novel objects and adapt to variations in shape, compliance, or surface texture. Recent work  [252, 271, 272, 273] showed that tactile-informed IL can substantially improve grasp stability, reduce failure rates under visual occlusion, and enable fine in-hand reorientation without explicit object models.
Overall, tactile sensors are becoming an indispensable component of dexterous robotic end-effectors, bridging the gap between perception and control. Future research is expected to focus on developing scalable, durable, and multimodal tactile arrays, as well as fusing tactile signals with vision and proprioception to support generalizable manipulation policies in open-world environments.
IV-E Impact of End-effector Design on Imitation Learning Performance
While prior research has primarily focused on algorithmic innovations, the influence of end-effector morphology, actuation, and sensor configuration on IL performance has received comparatively limited attention. These factors can significantly affect data efficiency, policy generalization, and task success rate through the following mechanisms.
IV-E1 Morphology:
Redundant Degrees of Freedom and Ergonomic Alignment. High-DoF hands (e.g., Shadow Hand, 20+ DoF) can replicate fine human manipulations such as fingertip control, but they increase the dimensionality of the action space, leading to severe sparsity issues in IL. The RH20T dataset [139] shows that the demonstration data requirement for high-DoF hands grows exponentially. In contrast, three-fingered grippers (e.g., DoraHand [235]) trade off some dexterity for reduced policy learning complexity, achieving better zero-shot generalization in cross-domain tasks such as those in BridgeData V2 [274]. Anthropomorphic hands (e.g., ILDA Hand [148]) with geometric consistency to the human demonstrator’s workspace reduce cross-domain mapping errors. Antotsiou et al. [23] highlighted that differences in morphology, degrees of freedom, and joint constraints between the human hand and the robot hand can amplify retargeting errors and consequently raise the likelihood of task failure.
IV-E2 Actuation:
Trade-Offs Between Transmission Dynamics and Controllability. Hands such as the Shadow Hand use tendon-driven actuation, where compliance introduces nonlinear friction and hysteresis, increasing noise in the action–state mapping of demonstration data. Tendon-driven systems introduce nonlinearities such as friction and hysteresis in the transmission chain, which are known to degrade control fidelity in dexterous manipulators [187]. Rigid linkage actuation (e.g., BarrettHand [222]) reduces dynamic errors, but underactuated designs (e.g., adaptive grasping) limit the reproduction of fine finger motions. Direct-drive hands (e.g., LEAP Hand [209]) achieve millimeter-level control precision through joint-level motors but suffer from increased inertia due to motor mass, requiring IL policies to learn dynamic feedforward compensation.
IV-E3 Sensor Configuration:
Complementarity in Multimodal Perception. High-resolution tactile sensing (e.g., GelSight  [260]) can capture micro-force adjustments and slip signals in human demonstrations, significantly improving IL’s understanding of contact dynamics. Huang et al. [252] reported that removing tactile input causes task success rates to plummet to near-random levels, underscoring the critical importance of tactile information for successful multi-object in-hand manipulation. Differences in sampling frequency (e.g., vision at 30 Hz vs. tactile sensing at 1 kHz) introduce synchronization challenges. Lin et al. [285] addressed this via a cross-modal alignment network, reducing motion consistency error by 25% in bimanual coordination tasks. NeuralFeels [275] aligns vision with tactile streams via global timestamps and fuses both modalities in a joint neural implicit field, cutting trajectory-consistency error by 25% for precise in-hand manipulation.
Future end-effector designs should follow a “task–morphology–algorithm” co-optimization paradigm. High-dexterity tasks (e.g., surgical suturing) favor fully actuated hands with dense tactile arrays (e.g., Shadow Dexterous Hand [150]) and employ hierarchical IL frameworks [74] to mitigate the curse of dimensionality. Low-cost generalization tasks (e.g., household tidying) use three-fingered grippers with vision-dominant sensing, combined with domain randomization [276] to improve cross-object robustness. Real-time collaboration tasks (e.g., human–robot co-manipulation) balance actuation latency and control precision. For instance, RAPID Hand [277] achieves <7<7 ms response delay through lightweight design.
TABLE VI: Comparison of Teleoperation Data Collection Approaches
Approach Key Characteristics Advantages Limitations
Vision-Based Systems [278, 18, 17, 78, 279, 20, 19, 280, 114, 281] Non-contact image-based motion capture Single or multi-camera configurations RGB or RGB-D inputs, frequently accompanied by hand pose estimation Low-cost hardware (especially single camera) Easy to deploy and user-friendly No wearables required Susceptible to occlusion and lighting conditions Limited 3D precision and depth estimation Latency due to vision processing
Mocap Gloves [282, 16, 283] Wearable gloves with IMUs, flex sensors, or magnetic sensors Real-time hand joint tracking Often used with external cameras or LiDAR High joint-level accuracy (millimeter-level) Fast response time Robust under varied visual conditions High equipment cost Requires calibration and setup Potential discomfort for extended use
VR/AR Controllers [14, 284, 285, 283, 286, 287, 288] Head-mounted display with spatial tracking Hand-held controllers with 6-DoF input Immersive virtual environments Intuitive and interactive control Multimodal feedback (e.g., visual and occasional haptic) Consumer-grade accessibility Dependence on virtual environments Challenging deployment and non-negligible control latency Limited haptic realism and force feedback
Exoskeleton and Bilateral Systems [289, 290, 291, 292, 77, 143, 293, 144] Direct joint-space mapping Includes bilateral force and torque feedback Leader-follower robot configurations High-fidelity control suitable for precision tasks Enables tactile interaction and real-time feedback Complex and bulky setups Expensive hardware and maintenance Restricted operator mobility
V Dexterous Manipulation with Teleoperation and Video Demonstration
This section discusses two main approaches to dexterous manipulation: acquiring demonstration data through teleoperation, and learning directly from demonstration videos. Teleoperation facilitates direct human control over robotic systems, leveraging human expertise for learning complex manipulation tasks, while video-based learning harnesses rich visual data to enable autonomous skill acquisition from natural human demonstrations. Together, these approaches address challenges in teaching robots fine-grained manipulation behaviors across diverse scenarios. The section further discusses datasets and benchmarks in supporting imitation learning frameworks that underpin these methodologies.
@@ V Dexterous Manipulation with Teleoperation and Video Demonstration
Learning dexterous manipulation skills from demonstration videos offers a promising alternative to teleoperation by leveraging abundant visual data to teach robots complex behaviors. Unlike direct control, video-based learning enables robots to observe human experts performing tasks in diverse and unstructured environments, capturing rich contextual and temporal information. Advances in computer vision and representation learning facilitate extracting meaningful features from raw videos, which can be mapped to robot actions through imitation learning frameworks. This approach not only reduces the reliance on specialized teleoperation hardware but also broadens the scalability of training data, ultimately enabling robots to acquire fine-grained manipulation capabilities from natural human demonstrations.
@@ V-A Teleoperation Systems for Dexterous Manipulation
To accurately convey human operators’ intentions to robotic systems, previous works have employed a wide range of human-robot interaction devices. Human operators with work experience can easily identify the current state of a robot through the image; however, accurately translating human instructions to robot actions remains a challenge. Some traditional controllers act on this: 1) joysticks [294] 2) haptic devices [295]; However, manipulation tasks often involve delicate movements and complex interactions, such as grasping, moving, and positioning small or irregularly shaped objects. These tasks necessitate devices that can offer dexterous interfaces to ensure the safety and efficacy of the robot’s actions. Precision and real-time feedback are crucial. Commonly used devices include: 1) cameras [18, 296, 17, 19, 280, 278, 279, 78, 20]; 2) mocap gloves [297, 298, 282, 299, 300, 301]; 3) VR/AR controllers [302, 303, 14, 304, 305, 78, 284, 27, 286, 306, 307, 308]; 4) exoskeletons and bilateral systems [289, 291, 292, 77, 143, 293]. A comparative summary of representative teleoperation approaches is presented in Tab. VI, highlighting their key characteristics, advantages, and limitations to facilitate effective human-robot interaction.
@@ V-A4 Exoskeleton and Bilateral Systems
Assessing teleoperation interfaces requires considering both interface fidelity and the quality of collected demonstrations. Key evaluation criteria include spatial and temporal tracking precision, the consistency between operator commands and robot execution under communication latency, control bandwidth, and the resolution of feedback. Demonstration quality is typically reflected in task success rate, trajectory smoothness, and the frequency of effective error recovery during operation. An important downstream measure is the interface-to-policy transfer efficiency, which captures how interface characteristics influence the performance and generalization of imitation learning policies. Interfaces with high fidelity and low latency generally enable richer demonstrations and yield superior policy performance in fine-grained manipulation tasks, while lower-bandwidth and less complex interfaces may be sufficient for coarse and high-level motion primitives. Incorporating these metrics together with empirical evidence provides a principled basis for interface comparison, supporting data collection strategies that balance hardware complexity, operator skill, and final policy effectiveness.
TABLE VII: Comparison of Video-based Imitation Learning Approaches
Method Category Representative Method Data Source Key Characteristics
Motion-centric Imitation Learning DexMV [309] Robotic Telekinesis [17] Track2Act [310] Unlabeled third-person real-world videos Unlabeled third-person unpaired videos Unlabeled videos, limited real-world data Reconstructs 3D hand-object trajectories or end-effector trajectories from videos and maps them to robot actions, enabling cross-domain imitation and improved generalization.
Synthetic Video for Policy Learning Gen2Act [311] NIL [312]
Text-generated unlabeled synthetic videos Diffusion-generated unlabeled visual videos
Uses language-conditioned video generation to synthesize training demonstrations, removing the need for expert data. Learns policies via perceptual similarity in generated videos, supporting large-scale training and generalization.
Representation Learning for Generalization Ag2Manip [313] Unlabeled human action videos Learns agent-agnostic action embeddings to support imitation across different robot embodiments and improve generalization.
Task-specific Architectures and Learning Objectives Bi-KVIL [314] Rank2Reward [315] ViViDex [316] Unlabeled third-person dual-hand videos Unlabeled videos Unlabeled noisy videos with no ground-truth state Specialized frameworks modeling bimanual coordination, reward learning without labels, and hierarchical learning architectures for robust imitation.
TABLE VIII: Categorization of Benchmark Datasets for Robotic Imitation Learning
Method Category Representative Datasets Scale Sensory Modalities
Human-teleoperated Datasets MIME [138] RH20T [139] BridgeData [140] BridgeData V2 [274] 8,260 demos, 20 tasks 110K+ multimodal sequences 7,200 demos, 71 tasks, 10 domains 60,096 trajectories, 24 environments RGB-D, Kinesthetic trajectories RGB-D, Audio, Tactile, Proprioception Egocentric vision, Action sequences Vision, Language embeddings
Augmented Datasets RoboAgent [317] 7,500 → 98K* augmented trajectories RGB, Semantic masks
Synthetic Datasets MimicGen [318] 200 → 50K synthesized demos, 18 tasks Object poses, Motion primitives
Dexterous and Bimanual Manipulation Datasets ARCTIC [319] DexGraspNet [320] OAKINK2 [321] 2.1M frames (3D meshes) 1.32M grasps, 5,355 objects 627 sequences, 4.01M frames Multi-view RGB, Contact dynamics Tactile sensing, Physics simulation Multi-view capture, 3D pose annotations
*RoboAgent employs semantic-preserving augmentation techniques to scale demonstrations
V-B Learning from Video Demonstrations for Dexterous Manipulation
Video-based demonstration learning offers a promising solution to the challenge of acquiring robotic manipulation skills in the absence of expensive expert demonstrations. A summary of representative approaches within this paradigm is presented in Tab. VII. By extracting structured behavioral cues from a wide range of human videos—including recorded, synthetic, and web-sourced data—these methods transform raw visual observations into robot-interpretable representations, thereby facilitating robust generalization across diverse tasks, objects, and embodiments.
V-B1 Motion-centric Imitation Learning
A major line of work adopts motion-centric imitation, where human hand–object interactions are converted into robot-executable trajectories. Qin et al. [309] propose a method that reconstructs 3D hand–object poses from third-person human videos and re-targets them to dexterous robot hardware, enabling high-fidelity policy learning without the need for direct teleoperation or egocentric demonstrations. Building upon this approach, Bharadhwaj et al. [310] predict object-relative end-effector trajectories from unstructured videos and fine-tune residual policies on robots, thereby achieving generalization across novel scenes and object configurations with minimal real-world training. Complementary to these approaches, Sivakumar et al. [17] directly map unstructured human hand motions from third-person videos to dexterous robotic hands, supporting real-time teleoperation without paired demonstrations.
V-B2 Synthetic Video-driven Imitation Learning
Another direction leverages synthetic video as training signals for policy learning. Bharadhwaj et al. [311] propose a method that generates realistic manipulation videos from textual prompts and internet instructions, which are then employed to train robot policies capable of generalizing to unseen categories and actions. Building upon this approach, Albaba et al. [312] introduce a diffusion-based video generation framework that replaces explicit action labels with perceptual similarity, providing a scalable pathway for training robots with diverse morphologies. By eliminating the reliance on real or expert demonstrations, these approaches significantly expand the scalability of video-based demonstration learning.
V-B3 Representation Learning for Generalization
A third stream of research focuses on representation learning for generalization, aiming to decouple policy learning from specific tasks and embodiments. Singh et al. [322] propose a method that spatially aligns human hands and objects in 3D, producing consistent motion patterns that can pretrain general-purpose manipulation policies and enable task-agnostic skill acquisition. Building upon this idea, Li et al. [313] learn shared agent-agnostic visual–action embeddings from diverse human operation videos, achieving robust sim-to-real transfer even on previously unseen tasks and object configurations. These approaches highlight the potential of abstracting beyond raw trajectory reconstruction to acquire transferable skills across domains.
TABLE IX: Rubric for Imitation Learning Dataset Quality Assessment
Criterion Description Example Indicators
Sensor Modality Richness Extent and variety of sensing channels provided. RGB, depth, tactile, proprioception, force–torque, audio
Annotation Quality Accuracy and consistency of labels across modalities and tasks. Object pose precision, end-effector position, task boundary clarity
Task & Scene Diversity Breadth of tasks, objects, and environmental variations covered. Number of object categories, scene layouts, lighting conditions
Physical Realism Degree to which simulated/augmented data approximates real-world physics and appearance. Visual fidelity, dynamics accuracy, sim-to-real transfer success
V-B4 Task-specific Architectures and Objectives
Finally, several task-specific architectures and learning objectives have been designed to address unique challenges in video-based imitation. Gao et al. [314] introduce a hybrid master–slave modeling strategy with low-level geometric constraints for dual-hand coordination, enabling robust reproduction of bimanual manipulation in cluttered environments. Building upon this idea, Shaw et al. [323] extract vision–action–force relations from human demonstrations to provide physical and semantic priors for dexterous robot learning, reducing the dependence on paired demonstrations. Extending this approach, Chen et al. [316] present a three-stage framework that combines video demonstrations, reinforcement learning-based trajectory optimization, and vision-only policy learning without privileged object information, thus enabling physically executable policies from noisy demonstrations. Yang et al. [315] address the reward specification bottleneck by constructing progress-based rewards from temporal rankings of video frames, supporting both reinforcement and adversarial imitation without access to action annotations. Collectively, these approaches advance the practicality of video-based demonstration learning by tackling challenges such as bimanual coordination, noisy inputs, reward learning, and task-specific constraints.
V-C Datasets and Benchmarks
V-C1 Human-teleoperated Datasets
A straightforward approach to collecting data for IL is through human teleoperation, where operators directly control robots to perform tasks. MIME [138] represents an early large-scale effort in this direction, providing 8,260 human–robot demonstrations across 20 diverse tasks, along with both human demonstration videos and kinesthetic robot trajectories. RH20T [139] extends this paradigm by scaling up to over 110,000 multimodal manipulation sequences and incorporating richer sensing modalities—visual, tactile, audio, and proprioceptive—captured through teleoperation interfaces with force–torque sensors and haptic feedback. This multimodal design enables one-shot IL across a wider range of tasks, robots, and environments. Building on the same teleoperation foundation, BridgeData [140] focuses on cross-environment generalization, offering 7,200 demonstrations across 71 tasks in 10 environments, primarily in kitchen settings. Its successor, BridgeData V2 [274] expands the dataset to 60,096 trajectories in 24 environments, covering tasks from pick-and-place to complex manipulations, and supporting scalable robot learning with multi-task and language-conditioned objectives.
V-C2 Augmented Datasets
To overcome the time-consuming and labor-intensive process of collecting large numbers of human demonstrations for IL, some datasets leverage data augmentation to expand existing demonstrations. RoboAgent [317] begins with 7,500 teleoperated trajectories and scales them up to roughly 98,000 using semantic augmentations, eliminating the need for additional human or robot effort. CyberDemo [324] adopts a similar strategy, collecting demonstrations in both simulated and real environments before applying extensive augmentations that introduce visual and physical variations, thereby improving policy robustness and generalization.
V-C3 Synthetic Datasets
Some datasets employ simulated or synthetic demonstration generation systems to greatly expand the amount of training data from a small set of human demonstrations. MimicGen [318] generates over 50,000 demonstrations across 18 tasks from roughly 200 human examples by adapting object-centric manipulation behaviors to new contexts via trajectory transformations. IntervenGen [325] builds on this idea by autonomously producing large sets of corrective interventions from minimal human input, thereby improving policy robustness to distribution shifts. DiffGen [326] adopts a similar strategy but incorporates differentiable physics simulation, photorealistic rendering, and vision–language models to create realistic robot demonstrations directly from text instructions.
V-C4 Dexterous and Bimanual Manipulation Datasets
Focusing on dexterous manipulation and hand–object interactions, ARCTIC [319] offers 2.1 million videos with precise 3D hand–object meshes and dynamic contact data, enabling the study of bimanual manipulation of articulated objects. DexGraspNet [320] provides 1.32 million grasps for 5,355 objects using ShadowHand, with each grasp physically validated in simulation to ensure stability. OAKINK2 [321] extends the scope to real-world bimanual object manipulation, comprising 627 sequences and 4.01 million frames from multi-view captures with detailed pose annotations for human bodies, hands, and objects. Tab. VIII presents a concise summary of the representative benchmarks and datasets discussed above.
From the perspective of dataset usability, several interdependent factors critically shape the practical value of an imitation learning dataset (as shown in Table IX). Foremost among these is sensor modality richness: datasets offering more comprehensive multi-modal signals (e.g., depth image, tactile feedback, proprioceptive states, force–torque readings, audio), generally deliver superior informational content and flexibility compared to RGB-only collections. Such multi-modal data facilitates more accurate perception, reasoning over physical contacts, and cross-modal policy learning. Annotation quality is another decisive factor, precise labels for object poses, manipulator or end effector positions, and task boundaries directly constrain the attainable performance of learned models. Task and scene diversity further influences generalization: datasets encompassing a wide spectrum of object categories, environmental configurations, and illumination conditions are better poised to support robust and transferable policies than those restricted to narrow, repetitive settings. For simulated or augmented datasets, physical realism is especially critical, as greater realism narrows the sim-to-real gap and boosts the likelihood that policies trained virtually will operate successfully on physical robots.